OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is perpendicular to y = -4x -2 and passes through the point (-16, -11).

OpenStudy (anonymous):

y+11=1/4(x+16) y+11=1/4x +4 y=1/4x -7

OpenStudy (anonymous):

y + y1 = m(x + x1) <<< what's the name of that equation again?

OpenStudy (anonymous):

point slope formula

OpenStudy (anonymous):

ah das right haha, thanks

OpenStudy (anonymous):

no problem